OpenStudy (anonymous):

Can someone explain the Cancellation process when multiplying fractions? Example: 23 3/4 x 95

OpenStudy (amistre64):

the cancellation process , if I understand you correctly, is simply the fact that a number divided by itself is equal to 1, and any number then multiplied by 1 is equal to itself.

OpenStudy (radar):

\[23+ 3/4 =95/4\] converting 23 3/4 to an improper fraction Did you understand the conversion of a mixed number to the improper fraction?

OpenStudy (amistre64):

take for example: 4 (5) ---- :5/5 = 1 5 (3) whats left is actually (4/3) times 1 = (4/3)

OpenStudy (radar):

In your example there is no "cancellation" simply multiply the numerator . Then multiply the denominator

OpenStudy (radar):

Lets say you had this fraction (4 x 2 x 3)/9 Cancelling is operating on a fraction without changing it value. Whatever you do to the bottom you must do to the top. In this case you can divide top and bottom by 3 getting (4 x 2)/3 This commonly called cancelling.

OpenStudy (anonymous):

Okay I converted 95 into an improper fraction and got 95/1

OpenStudy (anonymous):

no my quation is 23 3/4 x 95/1

OpenStudy (anonymous):

now my equation is 23 3/4 x 95/1, am sorry disregard that last message

OpenStudy (radar):

95/4 x 95/1=9025/4

OpenStudy (radar):

Are you saying you are not multiply the 23 3/4 ????

OpenStudy (radar):

Twenty three and 3/4 is equivalent to 95/4

OpenStudy (anonymous):

Yes I am trying to figure out whats the process, In my book it shows me how to do it but dont explain how to do it!!!

OpenStudy (anonymous):

This is what it say in the book: 23 3/4 x 95 solution: 95/1 x 3/4 = 285/4 = 71 1/4 95 x 23 = 2185 2185 + 71 3/4= 2256 1/4 2256 1/4 is the answer, but I need someone to explain this and break it don for me!!!

OpenStudy (radar):

O.K

OpenStudy (radar):

The book decided rather than convert 23 3/4 to a improper fraction they would do to separate actions and then combine. First they took the fractional part, the 3/4 and multiplied it by 95.\[3/4\times95=(3\times95)/4\] that valued is 285/4 was then converted to a mixed number and is obtained by dividing by 4 then placing the remainder was a 1 In other words 285/4 = 71 with 1 remaihder which is divided by 4 and placed as a fraction 1/4. Do you understand thus far?

OpenStudy (anonymous):

yes

OpenStudy (radar):

They then multiplied the 23 by 95 getting 2185. Now they add the two products that they got: 71 1/4 +2185. They simply added the whole numbert parts getting 2256 then adding the fraction 1/4 for the total of 2256 1/4.

OpenStudy (radar):

Now the answer I obtained earlier shown up above was 9025/4 If you divide the 4 into 9025 you get 2256 with a remainder of 1 or 1/4 Which is the same.

OpenStudy (radar):

This was not a good example and I did not see any accomplished.

OpenStudy (radar):

Good luck, and remember What ye doieth to the denominator yee must also do to the numerator.!!